### v1

# Label Set Optimization via Activation Distribution Kurtosis for Zero-shot Classification with Generative Models

Abstract
In-context learning (ICL) performance is known to be sensitive to the prompt design, yet the impact of class label options in zero-shot classification has been largely overlooked. This study presents the first comprehensive empirical study investigating how label option (e.g., lexical choice, order, and elaboration) influences zero-shot ICL classification performance. Our findings reveal that lexical choices for label names (e.g., agree vs.support in stance classification) play an important role, with effects also linked to label orders. An analysis of the model internal states further shows that optimal label names tend to activate fewer outlier neurons in the feed forward network. Based on this observation, we propose Label set Optimization via Activation Distribution kurtosiS (LOADS), a post-hoc approach requiring no gradient propagation. LOADS not only demonstrates effectiveness with only 100 unlabelled samples across different model types and sizes, but also shows cross-lingual transferability.
Label Set Optimization via Activation Distribution Kurtosis for Zero-shot Classification with Generative Models
Yue Li, Zhixue Zhao  and Carolina Scarton Department of Computer Science, University of Sheffield, UK {yli381,zhixue.zhao,c.scarton}@sheffield.ac.uk
1 Introduction
Nowadays, generative large language models (LLMs) are widely used directly on classification tasks with zero-shot ICL, where models are prompted to select an option from a pre-defined set of labels (Wang et al., 2022; Antypas et al., 2023; Gonen et al., 2023; Mu et al., 2024). While some classification tasks employ a relatively fixed set of lexicons to represent class labels – such as sentiment analysis (positive and negative) and textual entailment (entailment and contradiction) – other tasks may present more ambiguous choices in lexical selection. Stance classification, for instance, uses diverse pairs of antonyms to represent positive and negative stances across different datasets, e.g. agree-disagree and favor-against. As a result, when crafting prompts for these classification tasks, practitioners face decisions regarding label options in the prompt, such as the lexical selection and the ordering.
Figure 1: Illustration of the three aspects (i.e., lexical choice, label order and label elaboration) for designing the label option in the prompt in zero-shot ICL for classification, and our LOADS to post-hoc select the optimal label set (top half figure).
Despite studies suggesting the sensitivity of ICL to prompt design (Lu et al., 2022; Yoo et al., 2022; Wei et al., 2024; Mao et al., 2024; Liu et al., 2024a; Zhang et al., 2022; Liu et al., 2022a; Peng et al., 2024; Gonen et al., 2023; Mu et al., 2024), this subtle yet critical consideration of label options in prompt for zero-shot ICL has received limited attention. To fill in this research gap, we explore the impact of three types of label variants (i.e., lexical choice, label order and elaborations) in zero-shot ICL with both encoder-decoder and decoder-only LLMs. We ground our research on stance classification, a task where label adaptation is a known problem due to various label inventories in different studies (Hardalov et al., 2021). We demonstrate that the lexical choice of the label options significantly impacts model performance, also interacting with the label orders. Elaborating on task-related information (e.g. agree with the claim elaborating agree) has minimum effect.
Further analysis of the model internal state is also conducted to gain insights on optimal label selection. Inspired by recent studies on neuron analysis (Kuzmin et al., 2024; Voita et al., 2024; Stolfo et al., 2024), we investigate the neurons in the feed forward network (FFN) in the decoder of the LLMs. We empirically show that prompts with optimal label sets activate fewer outlier neurons. Consequently, we propose a new method, Label set Optimization via Activation Distribution kurtosiS (LOADS), to select optimal label sets for a given classification dataset in zero-shot ICL. LOADS could stably and effectively work with only 100 unlabelled samples of the validation dataset, also demonstrating cross-lingual transferability.
Our contributions can be summarised as follows:
  • The first benchmark on how variants of label options in prompts affect zero-shot ICL models’ performance for classification tasks. We provide useful recommendations on label designing to practitioners working on zero-shot classification with LLMs.
  • Empirical demonstration that zero-shot ICL performance negatively correlates with the number of outlier neurons in FFN when varying the lexical choices for label options. The correlation holds true across diverse English stance classification datasets with different models.
  • A novel and efficient post-hoc method (LOADS) for label selection in zero-shot ICL. Comparing with common strategies in practice, our approach demonstrates statistically significant performance improvements across different models types, model sizes and languages with only 100 unlabeled data samples.
Sections 2, 3 and 4 present our experimental setups, results and discussions for the impact of label options in zero-shot ICL. Then, we describe our neuron analysis of the lexical choice in label options in Section 5 and our proposed method for selecting optimal label sets in Section 6.
2 Prompting with Varied Label Options for Zero-shot Classification
In zero-shot ICL for classification, a common approach is to provide a set of class label options in the prompt in order to instruct the LLMs to choose one of the options as the classification prediction. Although the label option is a subtle component in the prompt, we are interested in whether it has significant impact on model performance.
Specifically, we explore three types of variants around label options in the prompt: (1) lexical choice; (2) label order; and (3) label elaboration. To properly measure the impact of these factors, we only manipulate with the label options within the same prompt template (Section 3). We show examples of the three variants in Figure 1.
Lexical Choice
We use single-word synonyms to represent class labels (e.g. support and agree), forming various label sets. For each dataset, we compare the zero-shot ICL performance when LLMs are prompted to select from different label sets, as illustrated in Figure 1. For this purpose, we design a pipeline to create a pool of label sets:
  1. 1.
    Collect a seed set of label names. We obtain this set by collecting the label names in the datasets we experiment with.
  2. 2.
    Expand label sets with WordNet and LLMs. We use WordNet (Fellbaum, 1998) as a reliable source and Claude [^1] https://claude.ai/new as a supplementary source to obtain synonyms for label names in the seed set. For pairs of label names with semantically opposite meanings (such as "agree" and "disagree"), we also consider antonyms to avoid potential ambiguity and present clear contrast for the predicted models.
  3. 3.
    Manual selection. We manually select the appropriate label sets generated by Claude to alleviate the impact of noisy label names.
The label names are arranged in the sequence presented in their original study (see Table 1). We refer to this arrangement as the default order.
Label Order
We consider every possible order of the single-word labels in the prompt and compare the model performance against that obtained with the default order. For binary datasets, there is only one alternative arrangement besides the default order, while $N$-way multi-class classification would yield $N!-1$ alternative orders.
Label Elaboration
We investigate whether transforming single-word labels (e.g., "agree") into more detailed phrases (e.g., "agree with the claim") has an impact on model performance. On the one hand, elaborating on task details with the label may provide the model with a stronger alignment signal between the label and the task, emphasizing what the label is referring to. On the other hand, it also increases the label length and may introduce noise in the prompt (Liu et al., 2024a). Therefore, we design three levels of elaborations (shorted for E1, E2 and E3) by progressively adding more task-related (and potentially redundant) information to the single-word label names, as shown in Figure 1.
3 Experimental Setups
Datasets
We focus on the stance classification task due to its rich label inventories across various readily available datasets. We consider four binary and five multi-class datasets from existing English stance classification benchmarks (Schiller et al., 2021; Hardalov et al., 2021; Chen et al., 2023), as shown in Table 1. The nine datasets cover different domains, such as social media posts, news articles and online debates forums. The target of the stance varies from non-phrase topics (e.g., "immigration" or "school uniform") to phrases or statements (e.g., news headlines, claims or rumourous tweets). Details can be found in Appendix A.
Dataset Name Labels
scd
(Hasan and Ng, 2013)
original: for, against
best: pro, con
perspectrum
(Chen et al., 2019)
original: support, undermine
best: validate, refute
snopes
(Hanselowski et al., 2019)
original: agree, refute
best: affirm, refute
ibmcs
(Bar-Haim et al., 2017)
original: pro, con
best: endorse, deny
vast
(Allaway and McKeown, 2020)
original: pro, con, neutral
best: confirm, dispute, neither
emergent
(Ferreira and Vlachos, 2016)
original: for, against, observing
best: endorse, reject, neutral
semeval
(Mohammad et al., 2016)
original: favour, against, neither
best: accept, reject, neutral
rumoureval
(Gorrell et al., 2019)
original: support, deny, query, comment
best: confirm, reject, question, neutral
arc
(Hanselowski et al., 2018)
original: agree, disagree, discuss, unrelated
best: affirm, refute, discuss, unrelated
Table 1: Lists of the English stance classification datasets, their labels in original dataset, and the best labels on Flan-T5-xl with zero-shot ICL as an example to justify our motivation on LOADS.
Models
We cover both encoder-decoder and decoder-only LLMs, and experiment with the prevalent open-sourced Flan-T5 (Chung et al., 2024), Llama 3 and Llama 3.1 (Dubey et al., 2024) model families as representatives for these two types of LLMs. We choose their moderate-sized instruction-tuned versions, Flan-T5-xl (3b), Llama -3-Instruct (8b) and Llama-3.1-Instruct (8b), as our primary models for investigation due to our hardware resources constraints and their decent zero-shot ICL performance (Aiyappa et al., 2024; Chung et al., 2024; Dubey et al., 2024).
$F_{score}$ perspectrum ibmcs snopes scd emergent semeval rumoureval arc vast
Llama 3 (8b) max 0.869 0.834 0.770 0.759 0.740 0.688 0.659 0.453 0.382
min 0.738 0.592 0.540 0.639 0.387 0.544 0.286 0.249 0.173
Original 0.799 0.679 0.656 0.709 0.467 0.643 0.521 0.364 0.219
avg±var 0.824±0.001 0.756±0.003 0.666±0.003 0.713±0.001 0.547±0.006 0.624±0.001 0.514±0.008 0.329±0.002 0.279±0.002
Llama 3.1 (8b) max 0.909 0.898 0.748 0.769 0.660 0.735 0.584 0.466 0.409
min 0.376 0.300 0.435 0.629 0.166 0.456 0.320 0.252 0.175
Original 0.809 0.759 0.660 0.749 0.494 0.676 0.480 0.426 0.237
avg±var 0.832±0.009 0.805±0.011 0.668±0.003 0.737±0.001 0.496±0.013 0.652±0.003 0.488±0.003 0.387±0.002 0.282±0.002
Flan-T5-xl (3b) max 0.940 0.960 0.809 0.776 0.743 0.706 0.761 0.685 0.496
min 0.836 0.807 0.576 0.449 0.487 0.166 0.281 0.358 0.155
Original 0.939 0.939 0.746 0.631 0.649 0.467 0.381 0.493 0.311
avg±var 0.899±0.001 0.901±0.002 0.695±0.004 0.661±0.009 0.626±0.003 0.539±0.012 0.520±0.010 0.507±0.008 0.328±0.008
Table 2: The maximum (max), minimum (min), average (avg), variance (var) of the model performance across different label names in the prompt for each validation set. The performance of the original label set (Original) is also included, showing that they fail to reach the maximum performance LLMs could get. The extent of the gap between the maximum and minimum performances is represented using colors: max--min$>0.3$, $0.2<$max--min$<0.3$. The greater the variance, the greater the impact of label lexical choice, and the greater the potential utility of optimizing the label set.
Prompt Template
We refer to the prompt template used in the supervised fine-tuning of Flan-T5 and Llama (Wang et al., 2022; Chung et al., 2024; Dubey et al., 2024). We fine-tune the template on validation sets to ensure the LLMs could follow the instructions on more than 95% of the samples. We also test other templates and find the results are consistent (e.g., prompting with label explanations in Appendix E). We present the results with the following prompt template in this paper: Given a [text1_name] and a [text2_name], detect the stance that the [text2_name] has towards the [text1_name]. There are {N} options: "{label0}", "{label1}, … , and {labelN-1}". Now complete the following example. [text1_name]: {text1}. [text2_name]: {text2}.
Evaluation
[…]
Implementation Details
To ensure reproducibility, we use greedy search for decoding [^3] We present the potential impact of the decoding strategy in Appendix D.. In more than 95% cases, LLMs exactly follow the instruction and output the stance name within the required stance options. Therefore, we directly use the model generation as the predicted label without post-processing or mapping. We run experiments on the validation set of each dataset. See Appendix C for details on the label sets experimented with. We exclude the topic/entities-based (such as stance towards Obama) stance classification datasets (i.e. scd, semeval and vast) in label elaboration experiments to avoid unnecessary ambiguity or bias during elaboration (e.g., the text to be classified may refer to anything from policy to personal behavior about Obama, and elaborating agree to agree with the opinion of Obama could lead to biased prediction).
4 Results and Analysis
We first discuss the impact of lexical choice, label order, and label elaboration on zero-shot ICL for classification (Subsection 4.1). Then we analyze whether optimal or poor label options could be generalized across datasets and models (Subsection 4.2). Suggestions for practitioners in zero-shot ICL for classification appear in Subsection 4.3.
4.1 Results
(a) top-k optimal label sets
(b) top-k poor label sets
Figure 2: The maximum performance gain (positive value) and drop (negative value) on each validation set after re-ordering the label names for the top-k optimal (a) and poor (b) label sets with Llama 3. Results on other two LLMs can be found in Appendix F.
Lexical Choice
As shown in Table 2, performance varies across datasets and models solely due to changes in the label names within the prompt. The variations are more pronounced than those reported in previous studies Mu et al. (2024). The gap between the highest and lowest performance exceeds 0.1 for all datasets and models, and surpasses 0.2 on more than half of the datasets. We observe that certain stance labels could potentially trigger biased predictions, leading to extremely low performance. For example, Flan-T5 tends to always output support when using support, deny, neither for the semeval dataset, and Llama 3 overly predicts con when prompted by the label set pro, con, neutral for the emergent dataset.
Label Order
Overall, we observe limited influence of label order (see Appendix F). However, we are also interested in extremes of performance change caused by shifting label orders, particularly the maximum performance gain and drop, and whether the extreme gain or drop correlates with the lexicons used for the label names. Therefore, for each dataset, we first select the top-k [^4] Due to the exponential increasing of label order options and our limited computational resources, we set k=15 for binary classification; k=10 for three-way classification and k=2 for four-way classification optimal and poor label sets based on their performance in the same order (i.e. the default order). We then examine the maximum increase or decrease of the performance after re-arranging the label orders for optimal and poor label sets, respectively. As shown in Figure 2(a), altering the order for the optimal label sets has the risk of high performance drops (e.g., even more than 0.2 on the binary classification snopes dataset). While performance improvements are possible, the gains are relatively limited (lower than 0.1 on all datasets). Conversely, re-arranging the order for the poorly performing label sets offers potential for substantial improvement (Figure 2(b)). This high improvements for certain label sets after re-ordering suggests that the poor performance may partly stem from the initial sub-optimal ordering.
Label Elaborations
Similarly, we select the top-k555k=15 for binary classification, k=30 for three/four-way classification. optimal and poor single-word label sets based on their performance, and examine the performance change after elaborations. As shown in Table 3, the models are robust to the elaborations for either optimal or poor single-name label sets, indicating that adding task related details or increasing the label token lengths brings limited impact on average. However, we also observe relatively large performance change on certain datasets. Specifically, for the rumoureval dataset with Llama 3, we notice a performance drop larger than 0.2 when elaborating an optimal label set. Performance on the ibmcs dataset with Llama 3.1 could increase 0.2 when elaborating a poorly-performing label set.
Dataset $E_{1}$ $E_{2}$ $E_{3}$
Opt. Sub-opt. Opt. Sub-opt. Opt. Sub-opt.
perspectrum 0.016 0.018 0.010 0.015 0.017 0.009
ibmcs 0.027 0.041 0.024 0.014 0.029 0.044
snopes 0.055 0.040 0.054 0.026 0.029 0.018
emergent 0.051 0.053 0.047 0.038 0.022 0.040
rumoureval 0.095 0.036 0.084 0.020 0.058 0.102
arc 0.017 0.024 0.015 0.021 0.035 0.049
Table 3: The average absolute performance change after elaborating for optimal or poor single-word label sets with Llama 3 ($E_{1}$, $E_{2}$, $E_{3}$ see Figure 1). Results of other LLMs could be found in Appendix G.
4.2 Analysis
A natural question arising from the observed substantial impact of the lexical choices for label options in zero-shot ICL is: Can optimal or poor label names be generalised across datasets or even models? If so, for any new datasets/models, we could potentially adopt the optimal label names we found on related dataset/models. Therefore, for each dataset $D_{i}$ on each LLM $M_{j}$, we select the top-$k$ label sets with optimal ($L^{optimal}_{D_{j}\_M_{j}}$) or poor ($L^{poor}_{D_{i}\_M_{j}}$) performance, and calculate the number of overlap between the label sets. We only focus on the positive and negative stances in this section to enable comparison across binary, three-way and four-way stance classification datasets.
Generalisation across datasets within the same model.
To analyse whether the optimal/poor label sets found on one dataset $D_{i}$ could be adapted to another similar dataset $D_{j}$ on the same LLM $M$, we calculate the overlap of top-k optimal/poor ($L_{D_{i}\_M}\bigcap L_{D_{j}\_M}$) between each dataset.
Our analysis reveals that there is no guarantee of generalizability for optimal/poor label sets across stance datasets on the same model. The level and type of generalization across datasets are also model-dependent. For Llama 3.1, there is always an overlap among the worst top-5 label sets across all the datasets, but not among the best top-5. That is, for all datasets tested, there are shared label sets that always result in sub-optimal performance when using Llama 3.1. Also, using supportive and opposed as label options leads to the worst performance on seven out of nine datasets – Llama 3.1 tends to overly generate supportive. However, no single label set consistently yields the best performance across more than half of the datasets with Llama 3.1. Flan-T5 follows a similar trend.
Conversally, for Llama 3, the level of transferability of the optimal and poor label sets across datasets is relatively similar to each other. Certain datasets, e.g. semeval, exhibit low transferability of optimal and poor label sets from other datasets (it overlaps with only four/three datasets considering top-5 label sets), while for vast, its top-5 optimal or poor label sets overlap with every dataset we tested when using Llama 3. Overall, the level of transferability for optimal label sets on Llama 3 is higher than Llama 3.1 and Flan-T5.
Generalisation across models with the same dataset.
To analyse whether the optimal/poor label sets found for dataset $D$ on model $M_{i}$ could be adapted to another model $M_{j}$, we calculate the overlap of top-k optimal/poor ($L_{D\_M_{i}}\bigcap L_{D\_M_{j}}$) between each model.
We find that optimal or poorly-performing label sets identified for Llama 3 on a specific dataset are more likely to transfer effectively to Llama 3.1 and Flan-T5, compared to transfer from Llama 3.1 to Flan-T5. However, the degree of generalisation varies depending on the dataset. We include the full statistics in Appendix H.
perspectrum ibmcs snopes scd emergent semeval rumoureval arc vast
Llama 3 (8b) -0.4921* -0.3787* -0.4359* -0.4217* -0.5781 -0.2618* -0.3764* -0.1662 -0.3639*
Llama 3.1 (8b) -0.4103* -0.3642* -0.1874 -0.0686 -0.4310* -0.2232* -0.3944* -0.1708* -0.1208
Flan-T5-xl (3b) -0.4476* -0.3638* -0.6014* 0.2353 -0.1089 -0.2881* -0.1714* -0.5742 0.1003
Table 4: Spearman correlation co-efficiency between model performance on validation set and kurtosis of neurons in the last layer. Mark with * when p value is lower than 0.05.
4.3 Suggestions to Practitioners
Based on our results and analysis, we provide the following suggestions to practitioners in zero-shot ICL for classification:
  1. 1.
    Lexical designing for the label names should be considered as an important step in prompt engineering.
  2. 2.
    Single-word class label without elaboration on task information is able to achieve high performance in most cases, i.e. adding extra information does not yield better results.
  3. 3.
    The optimal lexical choices for label options is both dataset-dependent and model-dependent.
  4. 4.
    If the practitioner has selected a set of optimal lexicons for label options based on a specific order, exploring alternative label orders can be redundant due to the limited performance gain brought from high computational costs. However, if a label set is chosen randomly, experimenting with different label orders may yield meaningful improvements (see Figure 2).
5 Neuron Analysis for Label Selection
Although our findings indicate the importance of label name selection for classification in zero-shot ICL, current studies lack consideration of this factor. Therefore, we conduct analysis to gain insights into the underlying mechanism of lexical choice for single-word label names. Inspired by studies on pre-trained language models or LLMs that indicate the correlation between neuron activation pattern in FFN and model performance (Kuzmin et al., 2023; Tang et al., 2024; Stolfo et al., 2024; Wu et al., 2024), we hypothesise the model performance influenced by label names may correlate with the number of outliers in the neurons within FFN in the decoder of the LLMs.
Methodology
For each FFN module in layer $i$ in the decoder, it can be denoted as follows:
$h^{i}=(\text{act\_fn}(\tilde{h}^{i}W^{i}_{1})\otimes\tilde{h}^{i}W^{i}_{3})\cdot W^{i}_{2}.$ (1)
where $\tilde{h}^{i}$ is the output hidden states from multi-head self-attention module. The activation function (act_fn) for Flan-T5 and Llama 3/3.1 is Gaussian Error Linear Unit (GELU) (Hendrycks and Gimpel, 2016) and Sigmoid Linear Unit (SiLU) (Hendrycks and Gimpel, 2016; Elfwing et al., 2018), respectively.
[…]
$\text{Kurtosis}[N_{I}]=\frac{\mathbb{E}[(N_{I}-\mu)^{4}]}{(\mathbb{E}[(N_{I}-\mu)^{2}])^{2}}$ (2)
where $\mu$ is the mean of $N_{I}$. For each dataset, we average the kurtosis score over the validation set for each candidate label set. We then calculate the Spearman correlation between model performance and the averaged kurtosis score.
Results
Figure 3: Macro-F1 vs. neuron activation kurtosis for the emergent dataset with different label options in the prompt for Llama 3.
Table 4 shows that, for most datasets, there is statistically significant negative correlation between model performances and kurtosis scores across models and activation functions, indicating that fewer outliers in the neurons of the final layers are associated with enhanced zero-shot ICL performance. We visualize an example of such correlation in Figure 3. This observation implies that the kurtosis score of neuron activation distribution in FFN of the last decoder layer of LLMs could potentially serve as an effective signal for selecting optimal label names in zero-shot classification.
perspectrum ibmcs snopes scd emergent semeval rumoureval arc vast
Llama 3 (8b) Lowest Kurtosis 0.8431 0.7684 0.5516 0.6698 0.5870 0.6445 0.4097 0.3662 0.3190
Original Label 0.8187 0.5485 0.4387 0.6504 0.3416 0.5565 0.3487 0.3130 0.2395
Llama 3.1 (8b) Lowest Kurtosis 0.8789 0.8856 0.6212 0.7274 0.6403 0.6581 0.3642 0.3637 0.2458
Original Label 0.8064 0.6783 0.4426 0.6983 0.4337 0.6492 0.3528 0.3784 0.2126
Gemma 2 (9b) Lowest Kurtosis 0.9110 0.9247 0.6233 0.7595 0.5989 0.6784 0.4896 0.4858 0.3439
Original Label 0.8966 0.8728 0.6454 0.7707 0.5704 0.6874 0.4778 0.4858 0.3383
Flan-T5-xl (3b) Lowest Kurtosis 0.9334 0.9380 0.6881 0.5997 0.5813 0.4951 0.4759 0.4688 0.3857
Original Label 0.9305 0.8971 0.7267 0.6341 0.5580 0.5697 0.3837 0.4628 0.3473
Flan-T5-xxl (12b) Lowest Kurtosis 0.9428 0.9644 0.6905 0.6158 0.5938 0.5257 0.3852 0.6151 0.4218
Original Label 0.9407 0.9630 0.7814 0.7598 0.5614 0.5697 0.2016 0.6065 0.3278
Table 5: Comparison of zero-shot ICL performance on test sets between prompting with LOADS-selected label names versus original dataset label names. We underline the highest model performance (statistically significant with paired chi-squared test).
French Portuguese
Llama 3 (8b) Lowest Kurtosis 0.5020 0.4528
Original Label 0.4544 0.4284
Llama 3.1 (8b) Lowest Kurtosis 0.4728 0.4278
Original Label 0.3731 0.3679
Flan-T5-xl (3b) Lowest Kurtosis 0.5137 0.3912
Original Label 0.4189 0.3317
Table 6: Performance when LLMs are prompted with English instructions (including label options) and French/Portuguese inputs. Label options are selected by LOADS with English validation data.
6 LOADS: Label set Optimization via Activation Distribution kurtosiS
Motivated by the above observation that the fluctuated zero-shot ICL performance caused by different label names could be attributed to the number of outliers in neurons in the last layer of LLMs, we propose LOADS to obtain an optimal label set for a given classification task in a post-hoc setting.
6.1 Method
We design a three-step pipeline based on LOADS for automatic label selection in zero-shot ICL:
  1. 1.
    Create a list of candidate label sets for class options in the prompt (see Section 2). The label names in each set should follow the same order.
  2. 2.
    Rank the list of label options based on the kurtosis score of the neuron activation in FFN of the last decoder layer (averaged across the validation set).
  3. 3.
    Choose the label set that leads to the lowest averaged kurtosis score.
6.2 Results
Evaluation
For evaluation, we compare the model performance of using label names selected by LOADS with the performance of using original label names from the dataset (i.e., the label names in Table 1), as using the original label name is the conventional and common practice when prompting the LLMs for classification tasks.
We randomly sample 100 data points from the validation set for label selection and test the selected label sets on the official test sets. Further, to examine whether our approach could generalize to other LLMs and LLMs with different model sizes, we also experiment with instruction-tuned Gemma-2-it (9b)  [^6] https://huggingface.co/google/gemma-2-9b-it and a larger version of Flan-T5, Flan-T5-xxl with 13 billion parameters.
Table 5 demonstrates that, on most of the datasets, employing LOADS to select label sets for zero-shot ICL prompts yields superior performance compared to directly using the original dataset label names. The improvement is consistent across diverse model architectures and sizes. [^7] We also test the prompt sensitivity in Appendix I.
Cross-lingual Transferability
Previous research (Zhang et al., 2023a) has demonstrated that for non-English datasets, using English task instructions (including label names) while keeping the input in the original language often yields superior performance than non-English instructions. Therefore, we investigate whether the optimal English label sets selected by LOADS on English dataset can also enhance performance on non-English datasets in this scenario.
We manually translate the English rumoureval Twitter test set into French and Portuguese. We select the optimal label set based on LOADS with 100 randomly sampled data from the English rumoureval validation set. We then prompt the LLMs with English task instructions and French/Portuguese inputs. Table 6 presents the results, indicating that the optimal English label set identified by LOADS also effectively improves performance on non-English datasets when the instruction is provided in English.
Data Efficiency of LOADS
To show the merit of data efficiency of LOADS, we randomly sample 50, 100, 300, 500 or 1000 data points from the validation set and compare the rankings of the label sets based on LOADS. Due to the resource restriction, we conduct experiments on snopes and emergent datasets with Flan-T5-xl and Llama 3.
The results show that the rankings of the top 5 label sets remain consistent across different sample sizes. It suggests that LOADS can achieve comparable performance even with a smaller number of unlabeled data samples than 100, further highlighting the data efficiency of our proposed method.
Alternative Methods
We explore other approaches to improve LOADS, including prompt perplexity (Gonen et al., 2023) (see Appendix J) and model internal representation of label names (Wang et al., 2023) (see Appendix K), but they did not yield any significant correlation with zero-shot ICL model performance.
7 Related Work
ICL performance
Most work on few-shot ICL focuses on cases where LLMs are directly prompted with $N$ demonstration examples. The studies highlight the substantial impact of example ordering (Lu et al., 2022), formatting (Yoo et al., 2022; Wei et al., 2024; Mao et al., 2024; Liu et al., 2024a), and examples selection (Zhang et al., 2022; Liu et al., 2022a; Peng et al., 2024). Some work focuses on improving few-shot ICL performance via the optimal selection or arrangement of examples (Liu et al., 2022a; Rubin et al., 2022; Lu et al., 2022; Zhang et al., 2022; Liu et al., 2024b; Xu et al., 2024), re-weighting examples (Yang et al., 2023), automatic reformat or generation of demonstration representations (Kim et al., 2022; Liu et al., 2023), and introduction of intermediate reasoning steps (Wei et al., 2022; Zhang et al., 2023b). However, the impact of lexical choices for label names in classification tasks has received little attention, with the only closely related work suggesting that models are likely to confuse class categories, given their similar key vectors in the attention modules (Wang et al., 2023). For zero-shot ICL, Mu et al. (2024) demonstrate the effect of using synonyms for class options, but they neither consider the order of the label options nor propose an effective strategy to choose the label names. Gonen et al. (2023) empirically show that perplexity could be an effective indicator for prompt selection, but they do not account for class options. Yet different model behaviors between few-shot ICL and zero-shot ICL have been frequently observed, suggesting that findings from few-shot ICL do not necessarily hold in the zero-shot context (Lin and Lee, 2024).
Prompt-Tuning and Verbalizer
Prompt-tuning aims to automatically find or generate an optimal discrete prompt (e.g., through gradient-based search (Shin et al., 2020; Shi et al., 2023) or fine-tuning (Gao et al., 2021; Le Scao and Rush, 2021; Deng et al., 2022)) or by training continuous prompt (Lester et al., 2021; Liu et al., 2022b). Verbalizer is commonly followed after prompt-tuning or pattern exploit training (Gao et al., 2021) for the languages models pre-trained with masked language modelling. It aims to map output predictions to labels for classification tasks, including manually created verbalizer (Schick and Schütze, 2021), search-based verbalizer that identifies label words automatically from the dataset (Gao et al., 2021; Shin et al., 2020), and soft verbalizers that uses continuous embeddings obtained through fine-tuning (Hambardzumyan et al., 2021; Cui et al., 2022).
8 Conclusion
We study the impact of label options in the prompt for classification in zero-shot ICL, including lexical choice, labels’ order, and label elaborations. We observe a significant effect of the lexicons used to represent label names in the prompt, also linking to the impact of the labels’ order. Through neuron activation analysis, we find that optimal label sets produce fewer outlier neurons in LLMs’ feed-forward networks. We then propose LOADS, a novel method for selecting optimal label sets using activation distribution kurtosis. Prompting with label set selected by LOADS consistently outperforms the use of original dataset labels across different models. Our approach is post-hoc, data-efficient and requires no gradient propagation or model fine-tuning. It also demonstrates cross-lingual transferability when using English instructions for non-English datasets. By showing that carefully selecting label sets based on neuron activation patterns can significantly enhance model performance without requiring additional training or labeled data, this paper has important implications for leveraging LLMs in zero-shot classification scenarios.
Limitation
Our experiments focused primarily on stance classification tasks, and future studies should explore whether these findings generalize to a broader range of classification tasks and domains. Also, although we examined multiple models from the Flan-T5 and Llama families, our study did not include other popular language models (such as Phi [^8] https://huggingface.co/microsoft/phi-2 or Mistral  [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2) due to computational resource limitation. Expanding the range of models would provide a more comprehensive understanding of label option’s impact across different architectures.
Another limitation of our study is English-language bias. Although we have explored the cross-lingual transferability to French and Portuguese on one dataset, more extensive multilingual testing is needed to ensure the approach’s effectiveness across diverse languages and cultures.
Our method, while more efficient than gradient-based approaches, still requires running inference on a subset of data to compute activation statistics. This may be challenging for resource-constrained environments or very large models. Also, we ensure the inclusion of samples for each class. The effectiveness of our method might vary when the validation set is highly imbalanced or even lack of data for the minority class. Also, the effectiveness of different distribution metrics is out of scope but we acknowledge that it may have significant improvement for our method.
Lastly, while we focused on technical performance, future work should consider potential biases introduced by label set choices and their implications for fairness and inclusivity in classification tasks. Addressing these limitations in future research will help to further validate and refine our approach to optimal label set selection for zero-shot ICL.
Ethics Statement
Our work has received approval from the Ethics Committee of our university and complies with X’s data policy for research.
Acknowledgement
This work is funded by EMIF managed by the Calouste Gulbenkian Foundation [^10] The sole responsibility for any content supported by the European Media and Information Fund lies with the author(s) and it may not necessarily reflect the positions of the EMIF and the Fund Partners, the Calouste Gulbenkian Foundation and the European University Institute. under the "Supporting Research into Media, Disinformation and Information Literacy Across Europe" call (ExU – project number: 291191).111111exuproject.sites.sheffield.ac.uk Yue Li is supported by a Sheffield–China Scholarships Council PhD Scholarship. The Portuguese and French translations of RumourEval were funded by the European Association for Machine Translation (EAMT). [^12] https://eamt.org We acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing.
Appendix A Datasets
We summarise the dataset characteristics in Table 7. For the datasets without official validation sets, we use the train/validation splits provided by Schiller et al. (2021).
Dataset Name Target Source Number of Label Sets
scd Topic Debates 31
perspectrum Claim Debates 31
snopes Claim News 31
ibmcs Topic Debates + Wikipedia 31
vast Topic Debates + Artificial 62
emergent Headline News 62
semeval Topic Social Media 93
rumoureval Tweet Social Media 248
arc Claim Debates 62
Table 7: Dataset characteristics (Hardalov et al., 2021) and the number of label sets we experiment with for each dataset.
Appendix B Data Leakage
As far as we know, Llama 3 and Llama 3.1 are not supervised fine-tuned with any public stance classification datasets. Flan-T5 is fine-tuned on Super-NaturalInstructions dataset (Wang et al., 2022), containing two English stance classification tasks (Kobbe et al., 2020) (i.e., task 209 and 513). The tasks are formed as a binary (“in favor” and “against”) and a three-way (“in favor”, “against” and “neutral”) classification, respectively. There is no overlapping between these two datasets and our nine experimented datasets.
Appendix C Label Pool Creation
Following the pipeline we described in Section 2, we collect the seed label sets from the nine stance classification datasets (see Table 1). For the semeval dataset, the label set in the original paper ("favor, against, neither" in Table 1) is slightly different from the set used in their published dataset ("favor, against, none"), so we consider both of them.
For positive and negative stance label names, we aim to acquire word-pairs with semantically opposite meaning. We first extract antonym for each positive and negative seed stance label from WordNet. Since we obtain limited antonyms in this way, Claude is then used to generate synonym for each seed positive-negative stance label pairs. An example of the prompt we used is: Provide 5 different pairs of synonyms for "support" and "deny". They are supposed to be labels for stance classification. We use WordNet to obtain synonyms for the rest of stance labels if there are any. For the label names that represent "neutral" stance in the original study, such as "observing" and "comment", we take "neutral" as their synonyms. Finally, we manually select the appropriate label names generated by Claude. The number of label sets we experiment with for each dataset is listed in Table 7.
Appendix D Decoding Strategies
We adjust the temperature (0.2, 0.4, 0.6, 0.8, 1.0, 1.2, 1.4) used for sampling-based decoding and compare their performances with the greedy search based performance for emergent and snopes dataset on Flan-T5-xl and Llama 3.
A temperature value larger than 1.0 – flattening the probability distribution – tends to harm the performance especially for Flan-T5-xl, which generates outputs irrelevant to stance. When temperature is lower than 1.0, introducing randomness in decoding through sampling may benefit the performance, but not significantly (in most of cases improvement is lower than 0.07). We summarise the maximum performance increase or decrease comparing with greedy search in Table 8.
Model Name Temperature snopes emergent
Flan-T5 0.2 0.021 0.049 0.037 0.049
0.4 0.055 0.066 0.066 0.068
0.6 0.033 0.099 0.034 0.064
0.8 0.033 0.145 0.066 0.107
1.0 0.022 0.189 0.036 0.149
1.2 0.037 0.264 0.046 0.196
1.4 0.015 0.428 0.042 0.361
Llama 3 0.2 0.050 0.063 0.068 0.035
0.4 0.044 0.047 0.073 0.073
0.6 0.040 0.062 0.075 0.059
0.8 0.027 0.062 0.066 0.107
1.0 0.054 0.086 0.100 0.103
1.2 0.038 0.072 0.069 0.171
1.4 0.030 0.104 0.128 0.142
Table 8: The maximum performance increase (+) and decrease (-) if adopting sampling-based decoding rather than greedy search.
Appendix E Prompting with Label Explanation
We investigate whether the performance variance caused by different lexical choices of the label names could be mitigated or lowered by including the explanation of the label names in the prompt. We experiment with emergent and snopes datasets on Flan-T5-xl and Llama 3. We add the following class explanations in the prompt template after the class options for snopes and emergent datasets respectively: (1) snopes: If the text supports that claim, answer with "{positive stance}"; if the text opposes the claim, answer with "{negative stance}"; (2) emergent: If the headline supports the claim, answer with "{positive stance}"; if the headline opposes the claim, answer with "{negative stance}"; if the claim is discussed in the headline but without assessment of its veracity, "{neutral stance}".
We observe that including these label name explanations in the prompt may help with the label sets that achieve the lowest zero-shot performance. As for the snopes dataset, its worst performance would increase from 0.5400 to 0.555 (Llama 3, labels: supportive and opposed) or from 0.5766 to 0.6568 (Flan-T5-xl, labels: for, against). As for emergent, its lowest performance would increase significantly from 0.3877 to 0.5600 with Llama 3 (labels: pro, con and neutral). However, when using Flan-T5-xl, the inclusion of the class explanation even decrease the worst performance from 0.4870 to 0.3775 (labels: support, deny and neutral).
More importantly, the benefits from label explanations in the prompt would not close the gap between the optimal and sub-optimal label sets, comparing the above improved performance with the maximum performances in Table 2.
Appendix F Label Order Results
We present the averaged absolute performance difference after re-ordering the label names in the prompt in Table 9. The influence is limited on average.
Dataset Flan-T5 Llama 3 Llama 3.1
perspectrum 0.0148 0.0372 0.0771
ibmcs 0.0195 0.0696 0.0961
snopes 0.0296 0.0772 0.1259
scd 0.0309 0.0288 0.0484
emergent 0.0211 0.0689 0.1578
semeval 0.0230 0.0304 0.0527
vast 0.0311 0.0465 0.0495
rumoureval 0.0355 0.0720 0.0439
arc 0.0152 0.0606 0.0612
Table 9: The average absolute performance change after re-ordering the label options in the prompt.
We supplement the maximum performance gain and drop on each dataset after re-ordering the label names for the top-k optimal and poor label sets with Llama 3.1 and Flan-T5-xl in Figure 4.
(a) Llama 3.1: top-k optimal label sets
(b) Llama 3.1: top-k poor label sets
(c) Flan-T5-xl: top-k optimal label sets
(d) Flan-T5-xl: top-k poor label sets
Figure 4: The maximum performance gain (positive value) and drop (negative value) on each dataset after re-ordering the label names for the top-k optimal and poor label sets with Llama 3.1 and Flan-T5-xl.
Appendix G Label Elaboration Results
We supplement the averaged absolute performance difference for each level of elaboration on Llama 3.1 and Flan-T5-xl in Table 10.
Dataset $E_{1}$ $E_{2}$ $E_{3}$
Opt. Sub-opt. Opt. Sub-opt. Opt. Sub-opt.
Llama 3.1 perspectrum 0.027 0.020 0.048 0.025 0.037 0.022
ibmcs 0.033 0.018 0.054 0.031 0.057 0.067
snopes 0.015 0.021 0.032 0.020 0.034 0.033
emergent 0.048 0.077 0.048 0.107 0.034 0.133
rumoureval 0.041 0.040 0.037 0.032 0.039 0.035
arc 0.032 0.046 0.029 0.036 0.042 0.046
Flan-T5-xl perspectrum 0.009 0.026 0.010 0.021 0.013 0.021
ibmcs 0.014 0.015 0.016 0.020 0.017 0.028
snopes 0.020 0.026 0.022 0.032 0.033 0.016
emergent 0.026 0.025 0.044 0.031 0.042 0.042
rumoureval 0.059 0.040 0.060 0.031 0.086 0.018
arc 0.034 0.042 0.025 0.038 0.031 0.041
Table 10: The average absolute performance change after elaborating for optimal or poor single-word label sets with Llama 3.1 and Flan-t5-xl ($E_{1}$, $E_{2}$, $E_{3}$ see Figure 1).
Appendix H Label Generalization Analysis
we report the overlap of top-5 optimal/poor ($L_{D_{i}\_M}\bigcap L_{D_{j}\_M}$) between each dataset on Llama 3, Llama 3.1 and Flan-T5-xl in Figure 5. The overlap of top-5 optimal/poor ($L_{D\_M_{i}}\bigcap L_{D\_M_{j}}$) between each model is presented in Figure 6.
(a) Llama 3 (Opt.)
(b) Llama 3 (Sub-Opt.)
(c) Llama 3.1 (Opt.)
(d) Llama 3.1 (Sub-Opt.)
(e) Flan-T5 (Opt.)
(f) Flan-T5 (Sub-Opt.)
Figure 5: Overlap among top-5 best (Opt.) or worst (Sub-Opt.) positive-negative stance label sets across datasets on Llama 3, Llama 3.1 and Flan-t5-xl respectively. Each entry $ij$ represents the number of overlapped best or worst label set between dataset on row $i$ and column $j$. The greater number indicate one label set is observed shared effectiveness across datasts.
(a) Best top5
(b) Worst top5
Figure 6: Overlap among top-5 best and worst positive-negative stance label sets on same dataset between different models. Each entry $ij$ represents the number of overlapped label sets for dataset $i$ (row $i$) between the models on column $j$. LLaMA3&LLaMA3.1: column $1$, LLaMA3&Flan-t5: column $2$, LLaMA3.1&Flan-t5: column $3$.
Appendix I Prompt Sensitivity Analysis of LOADS
To analyse the the prompt sensitivity of LOADS, we test it on different prompt templates, select the label set through LOADS, and then compare the performance with that on the label sets in the original dataset.
Due to the computational resource constraints, we manually craft two prompts and test LOADS with the four binary stance classification datasets on Llama 3. In the two prompts, we replace Given a [text1_name] and a [text2_name], detect the stance that the [text2_name] has towards the [text1_name] (see Section 3) with two different queries:
  1. 1.
    Prompt 1: What is the stance of [text2_name] towards [text1_name]?
  2. 2.
    Prompt 2: What stance does [text2_name] take regarding [text1_name]?
As shown in Table 11, although different prompts with the same label sets may result in performance changes as expected (compare with Table 2), LOADS is robust to different prompts used for the label selection. The performance gap between LOADS-selected and original label sets tends to be similar across prompts.
Dataset Lowest Kurtosis Original Label
Prompt 1 snopes 0.5926 0.4984
ibmcs 0.8737 0.7303
perspectrum 0.8925 0.8658
scd 0.6895 0.6860
Prompt 2 snopes 0.6191 0.5336
ibmcs 0.8619 0.7523
perspectrum 0.8921 0.8619
scd 0.6836 0.6931
Table 11: Performance comparison on Llama 3 when using LOADS-selected label sets (lowest kurtosis) and using original label sets (original label) with prompt 1 or prompt 2. The higher performance is underlined.
Appendix J Perplexity Analysis
As discussed in Section 7, Gonen et al. (2023) empirically show that zero-shot ICL performance is statistically negative correlated with the perplexity of the prompt with input. However, they did not take into account the label options in the prompt when calculating the perplexity. Therefore, we further investigate whether the perplexity is also correlated with the variance zero-shot ICL performance caused by different label names.
Specifically, we use the prompt template in Section 3 and calculate the perplexity of prompts with inputs and different label sets. Following Gonen et al. (2023), for each label set, we average the perplexity over the dataset. And then we adopt spearman correlation test between the averaged perplexity scores and model performances. Due to the computational restriction, we experiment with all the binary datasets on Flan-T5-xl and Llama3-8b. Since Flan-T5 is an encoder-decoder model where perplexity has a loose definition, we treat the encoder input as an empty string when calculating perplexity.
The results in Table 12 indicate that there is no statistically significant correlation between prompt perplexity and model performance if considering different label sets in the prompt.
perspectrum ibmcs snopes scd
Llama 3 coefficient 0.0068 0.1641 0.0394 0.1698
p value 0.9707 0.3774 0.8303 0.3608
Flan-T5 coefficient 0.0738 0.1733 -$0$0.2273
p value 0.6929 0.3511 0.7892 0.2187
Table 12: Spearman correlation between model performance and prompt perplexity. P-values are all larger than 0.05, indicating no statistical significance.
Appendix K Label Attention Key Similarity Analysis
In this section, we explore whether the closely related observation on few-shot ICL could be directly adopted to zero-shot ICL. Specifically, we focus on the study discussed in Section 7, where Wang et al. (2023) suggest that the when the LLM is prompted by demonstration with examples in a few-shot ICL setting, the model is likely to confuse the label categories if their key vectors in the attention modules are similar to each other.
Since this finding is easier to be tested on binary datasets, we experiment with the binary datasets on Llama 3 and Flan-T5-xl. We extract the key vectors in the attention module in each layer for each label name in the prompt. Then we calculate the cosine similarity between the vectors of two label names. Finally, we use spearman correlation test between similarity scores and model performances. As shown in Table 13, we do not observe statistically significant correlation between model performance and label key vector similarities.
perspectrum ibmcs snopes scd
Llama 3 coefficient -0.0181 -0.0051 -0.2791 -0.1696
p value 0.9244 0.9784 0.1283 0.3702
Flan-T5 coefficient -0.0595 0.0223 -0.0209 -0.0992
p value 0.7503 0.9047 0.9108 0.5953
Table 13: Spearman correlation between model performance and label’s key vector similarity.
Appendix L Layer-wise Output Projections Analysis
We hypothesize that LLM may jump to the output prediction at last layers when a sub-optimal label set is used in the prompt. Therefore, we extract the hidden states from each decoder layer and project them on the model vocabulary, so that we obtain the ranked position of the final predicted label’s token in each layer Elhage et al. (2021); Geva et al. (2022)..
we observe that the hypothesis indeed holds in certain cases. We show an example on rumoureval dataset where we compare the averaged rank of the correctly predicted label comment/neutral in each decoder layer when Flan-T5-xl is prompt to choose from support, deny, query, comment or support, deny, query, neutral. Label set support, deny, query, comment performs worse than the set endorse, deny, query, neutral on this dataset. As shown in Figure 7, when using the relatively optimal label set endorse, deny, query, neutral, the rank of the final predicted label tends to move closer to the top at an earlier stage.
Figure 7: The rank of the final correctly predicted label (comment orneutral) when Flan-t5-xl is prompted with two different label sets for rumoureval dataset.
Appendix M Human Translation Details
To translate the English Twitter rumoureval test set into French and Portuguese, we recruit volunteer students from translation studies in Brazilian and French universities. The students are given gift vouchers (0.6 pounds per tweet). Consent has been obtained from the students and our study has received approval from the Ethics Committee of our university.
We instruct the students to translate the tweets accurately, and preserve the original meaning, context, and tone of the tweet. They are also encouraged to leave notes for their translations. The translations are finished on Google Sheets.